# Semi-Supervised Semantic Segmentation with Professional and General Training

### v1

@@ Abstract
With the advancement of convolutional neural networks, semantic segmentation has achieved remarkable progress. The training of such networks heavily relies on image annotations, which are very expensive to obtain. Semi-supervised learning can utilize both labeled data and unlabeled data with the help of pseudo-labels. However, in many real-world scenarios where classes are imbalanced, majority classes often play a dominant role during training and the learning quality of minority classes can be undermined. To overcome this limitation, we propose a synergistic training framework, including a professional training module to enhance minority class learning and a general training module to learn more comprehensive semantic information. Based on a pixel selection strategy, they can iteratively learn from each other to reduce error accumulation and coupling. In addition, a dual contrastive learning with anchors is proposed to guarantee more distinct decision boundaries. In experiments, our framework demonstrates superior performance compared to state-of-the-art methods on benchmark datasets.
@@
Semi-Supervised Learning, Semantic Segmentation, Contrastive Learning
@@ 1 Introduction
Semantic segmentation, a critical task in computer vision, involves assigning a specific label to each pixel [1, 2, 3]. This pixel-level classification problem is notably more complex than image classification [4] or object detection [5] in the field of computer vision. Many approaches to semantic segmentation utilize an encoder-decoder architecture [6, 7], a strategy often employed in fields such as defect detection and autonomous driving. The dependence of supervised learning techniques on large-scale, accurately annotated datasets, which are often costly to produce, has sparked considerable interest in leveraging weakly labeled, easily labeled, and unlabeled data to boost model performance [8, 9, 10, 11, 12, 13]. Specifically, there has been significant progress in semi-supervised semantic segmentation, which combines a small amount of labeled data with a large volume of unlabeled data.
Figure 1: Illustration of some architectures. (a) Mean teacher [14, 15] with a teacher model $f(\bar{\theta})$ and a student model $f(\theta)$. (b) Cross confidence consistency [16] with two student models $f(\theta_{1})$ and $f(\theta_{2})$. (c) Cross pseudo supervision [11] with two student models $f(\theta_{1})$ and $f(\theta_{2})$. $x_{w}$, $x_{s}$ denote the samples with weak and strong augmentation, respectively.
@@
Semi-supervised semantic segmentation methods generally revolve around three core architectures: mean teacher [17, 14, 15], cross confidence consistency [16], and cross pseudo supervision [11], as illustrated in Fig. 1. In mean teacher architecture, the teacher model functions as an exponential moving average (EMA) of the student model [18], as depicted in Fig. 1(a). The other two architectures involve two student models with the same architecture learning from each other, as shown in Fig. 1(b) and Fig. 1(c). Additionally, two primary learning strategies are employed in semi-supervised semantic segmentation: consistency regularization [19, 20, 21] and entropy minimization [22, 23, 24]. Consistency regularization guarantees that a classifier generates similar predictions for a sample when subjected to various perturbations, and entropy minimization is a self-training process where a model is supervised using pseudo-labels generated by another model.
@@
However, these approaches do not pay attention to class imbalance issues, and most of them are likely to suffer from model coupling issues [18, 25], both of which can lead to more severe error accumulation. To overcome these limitations, we propose a Synergistic Training framework with Professional and General Training (STPG), as shown in Fig. 2. The simple yet effective framework consists of two modules: a professional training module and a general training module. The first module only selects pseudo-labels that are predicted consistently or highly mismatched between a professional student (Pro-Student) and a general teacher (Gen-Teacher), which can learn more minority class information and reduce error accumulation. In the second module, a general student (Gen-Student) is supervised by all pseudo-labels from a professional teacher (Pro-Teacher), which can learn more comprehensive semantic information. As the two modules are different, they are unlikely to be coupled. To our knowledge, this method represents a pioneering approach in semi-supervised semantic segmentation, addressing both class imbalance and model coupling issues simultaneously.
Figure 2: Illustration of the synergistic training framework. It contains a Gen-Student with a corresponding Gen-Teacher and a Pro-Student with a corresponding Pro-Teacher. Pro-Student acquires knowledge from Gen-Teacher to focus on high-quality and highly mismatched pseudo-labels and then transfers knowledge to Pro-Teacher. Gen-Student is supervised by all pseudo-labels from Pro-Teacher to obtain more comprehensive semantic information and then transfers knowledge to Gen-Teacher. Through loopy updates, high-quality and difficult samples can be continuously learned.
@@ 3.2 Synergistic Training
Professional Training Module. The traditional learning approach tends to focus on majority classes and is not able to sufficiently learn minority classes. Thus, we propose a pixel selection strategy for producing refined pseudo-labels to improve the performance. The loss is calculated using consistent and highly mismatched pseudo-labels from Gen-Teacher to supervise Pro-Student’s predictions instead of all the pseudo-labels.
@@
The highly mismatched pseudo-labels represented by $\hat{y}_{u}^{Hmis}$, as shown in Fig. 4(b), are computed in three steps.
@@
When it is large, this indicates a significant disagreement between the two models regarding Class $q$. This substantial disagreement suggests that Class $q$ is likely a minority class. The reason is that these classes are more prone to being misclassified as other classes, which results in fewer pixels being assigned to them in the pseudo-labels compared to other classes. We map the class-level mismatch scores onto the pixel predictions of both the Pro-Student and Gen-Teacher to generate two mismatch matrices, denoted as $\mathbf{I}_{(\cdot)}$. These matrices match the size of the image, with each element corresponding to the mismatch score of the predicted class, namely, $\mathbf{I}_{(\cdot)}[i,j]=I_{\mathbf{A}(\cdot)[i,j]}$. Here, $\mathbf{A}(\cdot)[i,j]$ represents the predicted class for the pixel located at the coordinates $(i,j)$ and $\cdot$ denotes $p_{u}^{Pro}$ or $\hat{p}_{u}^{Gen}$
@@ 4.1 Experimental Setting
Datasets. The Cityscapes dataset [44] is designed for semantic analysis of urban street scenes, includes 2975 training images with fine-grained annotations and 500 validation images, covering 19 classes. The PASCAL VOC 2012 dataset [45] serves as a standard benchmark for semantic segmentation with 20 object classes and 1 background class. It includes 1464 training images and 1449 validation images. Additionally, we use the SBD dataset [46] as an augmented set, which provides 9118 extra training images. Due to the coarse annotations in the SBD dataset, previous methods employ two primary partitioning strategies. The first strategy uses only the standard 1464 images as the entire labeled set, while the second strategy considers all 10582 images as potential labeled data. To ensure a fair comparison of our methods, we evaluate them using both a blender set (10582 potential labeled images) [15] and a classic set (1464 labeled images) [11].
@@ 5 Conclusion
In this paper, we propose a new framework for semi-supervised semantic segmentation. The STPG framework with professional and general training modules can reduce model coupling and error accumulation and retain more semantic information for minority classes. Within the framework, a dual contrastive learning can alleviate the influence of majority class dominance in the feature space. Our approach is easily adaptable to various semi-supervised semantic segmentation networks. The experimental results show that STPG surpasses the state-of-the-art techniques.

### v2

# Exploiting Minority Pseudo-Labels for Semi-Supervised Semantic Segmentation in Autonomous Driving

@@ Abstract
With the advancement of autonomous driving, semantic segmentation has achieved remarkable progress. The training of such networks heavily relies on image annotations, which are very expensive to obtain. Semi-supervised learning can utilize both labeled data and unlabeled data with the help of pseudo-labels. However, in many real-world scenarios where classes are imbalanced, majority classes often play a dominant role during training and the learning quality of minority classes can be undermined. To overcome this limitation, we propose a synergistic training framework, including a professional training module to enhance minority class learning and a general training module to learn more comprehensive semantic information. Based on a pixel selection strategy, they can iteratively learn from each other to reduce error accumulation and coupling. In addition, a dual contrastive learning with anchors is proposed to guarantee more distinct decision boundaries. In experiments, our framework demonstrates superior performance compared to state-of-the-art methods on benchmark datasets.
@@
Semi-supervised learning, Semantic segmentation, Autonomous driving
@@ 1 Introduction
Accurately segmenting different categories of road elements is crucial for effective semantic segmentation in intelligent transportation systems. Assigning a specific label to each pixel of pedestrians, vehicles, road signs, and other elements can enhance road safety and efficiency. The progress in this area can not only contribute to safer roadways but also pave the way for more intelligent and responsive driving systems. To this end, semantic segmentation becomes a critical task in computer vision, involves assigning a specific label to each pixel [1, 2, 3]. This pixel-level classification problem is notably more complex than image classification [4] or object detection [5] in the field of computer vision.
@@
Many approaches to semantic segmentation utilize an encoder-decoder architecture [6, 7], a strategy often employed in fields such as defect detection and autonomous driving. The dependence of supervised learning techniques on large-scale, accurately annotated datasets, which are often costly to produce, has sparked considerable interest in leveraging weakly labeled, easily labeled, and unlabeled data to boost model performance [8, 9, 10, 11, 12, 13]. Specifically, there has been significant progress in semi-supervised semantic segmentation, which combines a small amount of labeled data with a large volume of unlabeled data.
@@
Semi-supervised semantic segmentation methods generally revolve around three core architectures: mean teacher [17, 14, 15], cross confidence consistency [16], and cross pseudo supervision [11]. In mean teacher architecture, the teacher model functions as an exponential moving average (EMA) of the student model [18]. The other two architectures involve two student models with the same architecture learning from each other. Additionally, two primary learning strategies are employed in semi-supervised semantic segmentation: consistency regularization [19, 20, 21] and entropy minimization [22, 23, 24]. Consistency regularization guarantees that a classifier generates similar predictions for a sample when subjected to various perturbations, and entropy minimization is a self-training process where a model is supervised using pseudo-labels generated by another model.
@@
However, these approaches do not pay attention to class imbalance issues, and most of them are likely to suffer from model coupling issues [18, 25], both of which can lead to more severe error accumulation. To overcome these limitations, we propose a Synergistic Training framework with Professional and General Training (STPG). The simple yet effective framework consists of two modules: a professional training module and a general training module. The first module only selects pseudo-labels that are predicted consistently or highly mismatched between a professional student (Pro-Student) and a general teacher (Gen-Teacher), which can learn more minority class information and reduce error accumulation. In the second module, a general student (Gen-Student) is supervised by all pseudo-labels from a professional teacher (Pro-Teacher), which can learn more comprehensive semantic information. As the two modules are different, they are unlikely to be coupled. To our knowledge, this method represents a pioneering approach in semi-supervised semantic segmentation, addressing both class imbalance and model coupling issues simultaneously.
@@ 3.2 Synergistic Training
Professional Training Module. The traditional learning approach tends to focus on majority classes and is not able to sufficiently learn minority classes. Thus, we propose a pixel selection strategy for producing refined pseudo-labels to improve the performance. The loss is calculated using consistent and highly mismatched pseudo-labels from Gen-Teacher to supervise Pro-Student’s predictions instead of all the pseudo-labels. The minority-class pseudo-labels are approximated with highly mismatched pseudo-labels, based on the observation that minority classes are more likely to be misclassified as other classes.
@@
To identity mismatch pseudo-labels, we introduce a class-level mismatch score, which represents the proportion of a class being misclassified into other classes, based on the predictions of the Gen-teacher and Pro-student. This score reflects the degree of mismatch for a specific class. The highly mismatched pseudo-labels represented by $\hat{y}_{u}^{Hmis}$, as shown in Fig. 2(b), are computed in three steps.
@@
When the score is high, it indicates significant disagreement between the two modules regarding Class $q$. Such substantial disagreement suggests that Class $q$ is likely a minority class, as these classes tend to be misclassified more often, resulting in fewer pixels being assigned to them in the pseudo-labels compared to other classes.
We map the class-level mismatch scores onto the pixel predictions of both the Pro-Student and Gen-Teacher to generate two mismatch matrices, denoted as $\mathbf{I}_{(\cdot)}$. These matrices match the size of the image, with each element corresponding to the mismatch score of the predicted class, namely, $\mathbf{I}_{(\cdot)}[i,j]=I_{\mathbf{A}(\cdot)[i,j]}$. Here, $\mathbf{A}(\cdot)[i,j]$ represents the predicted class for the pixel located at the coordinates $(i,j)$ and $\cdot$ denotes $p_{u}^{Pro}$ or $\hat{p}_{u}^{Gen}$
@@ 4.1 Experimental Setting
Datasets. The Cityscapes dataset [44], designed for semantic analysis of urban street scenes, contains 2975 finely annotated training images and 500 validation images across 19 classes. However, the dataset suffers from a significant long-tail problem, with classes like roads and sky occupying the majority of pixels, while others like pedestrians and poles account for much fewer pixels, which severely hinders the model’s learning performance. In addition, due to the limitations of open source datasets on autonomous driving, we also verified the effectiveness of our approach on PASCAL VOC 2012 dataset which [45] serves as a standard benchmark for semantic segmentation with 20 object classes and 1 background class. And it includes 1464 training images and 1449 validation images. Additionally, we use the SBD dataset [46] as an augmented set, which provides 9118 extra training images. Due to the coarse annotations in the SBD dataset, previous methods employ two primary partitioning strategies. The first strategy uses only the standard 1464 images as the entire labeled set, while the second strategy considers all 10582 images as potential labeled data. To ensure a fair comparison of our methods, we evaluate them using both a blender set (10582 potential labeled images) [15] and a classic set (1464 labeled images) [11].
@@ 5 Conclusion
In this paper, we propose a new framework for semi-supervised semantic segmentation for autonomous driving. The STPG framework with professional and general training modules can reduce model coupling and error accumulation and retain more semantic information for minority classes. Within the framework, a dual contrastive learning can alleviate the influence of majority class dominance in the feature space. Our approach is easily adaptable to various semi-supervised semantic segmentation networks. The experimental results show that STPG surpasses the state-of-the-art techniques.